OpenStudy (anonymous):

substitution method x-3y=-11 3x+5y=-5

OpenStudy (solomonzelman):

x-3y=-11 ---> times -3 ---> -3x +9y =33 3x+5y=-5 ---> add this equation to it ----> 3x +5y =-5 ------------- 0x +14y = 28 14y=28 y=2 solve for x by plugging in y, x-3y=-11 x-3(2)=-11 x-6=-11 x-6+6=-11=6 x=-5 CHECK: (-5)-3(2)=-11 3(-5)+5(2)=-5 -5-6=-11 -15+10=-5 -11=-11 -5=-5 CORRECT: x=-5 y=2 or (-5,2)

OpenStudy (anonymous):

thank you :)

OpenStudy (solomonzelman):

Didn't see "by substitution method"

OpenStudy (amigatour):

x-3y=-11 ------> x=3y-11 substitute this to any of the two eqn systems. 3x+5y=-5 3(3y-11)+5y = -5 9y-33+5y = -5 9y+5y-33 = -5 14y -33 = -5 (opposite +33) 14y = 33-5 14y = 28 y= 2 Now, you do the rest...

OpenStudy (amigatour):

You plug the y=2 value in the system...

OpenStudy (amigatour):

to get the x-value(s)

OpenStudy (anonymous):

thank you amigatour

OpenStudy (amigatour):

yw.